SX AO commissioning

01:48 Get BCU 3 error on Power On. Stop after a 15 seconds.
  • Load programs with no problems
  • Set with no problems
mirrorctrl.L       |ERR|      1614|2016-10-18 01:48:34.697734| BCUCOMMANDHANDLE > BcuCommandHandler::handleBcuReply: REQUEST COMPLETED but FAILED :-(
mirrorctrl.L       |ERR|      1615|2016-10-18 01:48:34.697837| BCUCOMMANDHANDLE > Client timed-out: impossible to attach to the shared memory !
mirrorctrl.L       |ERR|      1616|2016-10-18 01:48:34.697862| BCUCOMMANDHANDLE > BcuCommandHandler::handleBcuRequest: REQUEST FAILED
mirrorctrl.L       |ERR|      1617|2016-10-18 01:48:36.700895|             MAIN > ******** FAILED PACKET FROM BCU 3 *******
mirrorctrl.L       |ERR|      1618|2016-10-18 01:48:38.202662|             MAIN > ******** FAILED PACKET FROM BCU 3 *******

02:10 Active Preset
  • Seeing, 1.3 to 2.0 on Guide image
  • 1.6 to 2.0 on DIMM

02:25 ACE Preset
  • RunAO fail. No communication to AdSec via fastlink
  • Call Lorenzo. They had changed fibers.
  • Go to Zenith so Lorenzo can reconfigure to FLAO setup.

02:38 Lorenzo done. Test comm. Good

02:40 ACE Preset
  • Bad seeing reconstructors
  • RunAO
  • Success
  • Seeing variable, from 1.0 to 2.2
  • Loop Open, skip frames

02:45 Switch to MOD1/2 for photometric standards and (if better seeing) transformation collection

03:05 Set DX AdSec with only a few BCU - 2 CA errors

07:17 Start AO work again. Seeing ~0.8 on DIMM

07:18 Check pointing at AO119

07:25 ACE preset
  • RunAO successful
  • Seeing getting worse, 1.5

07:43 ACE Preset
  • RunAO Fail with baysidex timeout
  • RunAO Fail with:
pyarg              |INF|       772|2016-10-18 07:43:01.895120|             MAIN > Setting tiptilt to amp= 0.00, freq= 0.00, offx= 0.00, offy= 0.00
pyarg              |ERR|       773|2016-10-18 07:43:01.933401|             MAIN > [Errno 4] Interrupted system call

07:45 ACE Preset
  • Failed
pyarg              |ERR|       808|2016-10-18 07:45:03.491387|             MAIN > Timeout waiting for slopecompctrl.L.STATUS (-5001) TIMEOUT_ERROR

07:47 Stop and start slope computer
  • force bin 1 on CCD39
  • OK

08:15 DX AdSec Failure. Get BCU 6 errro
  • Power off and adsc_stop/stat

08:17 Start dice-5 offsets
  • First luci image 0006.fits
  • Last luci image 0027.fits
  • 4 full dice 5
  • Error on offset
Tue Oct 18 08:33:44.070 2016 left Offset alert for LUCIFER: OffsetXY{FLAO}[142] - [AOException] RETRY: Stages position not reached (-5001) TIMEOUT_ERROR

08:39 Test one angle of NCPA amplitudes
  • Could not get AOS to show loop closed, so IIF could not perform offset
  • Preset to AO119
  • Seeing 1.8"
  • Current NCPA amplitudes definitely improve the image.
  • Seeing is bad enough that I can not tell if I can improve the image

09:25 ACQUIRE preset to OC_M34

09:35 Send LUCI preset
  • failed
Tue Oct 18 09:55:31.684 2016 left RunAO alert for LUCIFER: AcquireRef{FLAO}[169] - FLAO: unknown exception from arbitrator
Tue Oct 18 09:55:31.706 2016 left RunAO command failed from LUCIFER, type NORMAL, leftRePoint 0, rightRePoint 0
  • Resume loop at engineering GUI
  • press StartAO at AOSGUI to update status on AOS
  • iif_runao, TYPE=skipall to update TCS status
  • first image 0031.fits
  • 15,6 mag on pyrimid
  • 14.5 mag catalog
  • last image 0040.fits

10:25 Preset to dim off-axis reference start
Tue Oct 18 10:25:44.788 2016 left RunAO alert for LUCIFER: AcquireRef{FLAO}[217] - FLAO: unknown exception from arbitrator
Tue Oct 18 10:25:44.798 2016 left RunAO command failed from LUCIFER, type NORMAL, leftRePoint 0, rightRePoint 0
  • 15.8 mag on Pyramid
  • 14.5 mag catalog
  • images 0041.fits to 0045.fits

10:30 Optical loop data
  • 103046

10:33 Loop open, safeskip
Tue Oct 18 10:33:30.806 2016 left Offset alert for LUCIFER: Pause{FLAO}[51] - FLAO: unknown exception from arbitrator
Tue Oct 18 10:33:30.827 2016 left Offset command failed for LUCIFER, Sound=offsetFailed.wav

10:36 Preset to dim off-axis guide star
  • Failed, no star in TV image
Tue Oct 18 10:43:18.892 2016 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: PresetAO{FLAO}[247] - FLAO: unknown exception from arbitrator
Tue Oct 18 10:43:18.905 2016 left Preset command failed for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, offset, AO reference, Sound=presetFailed.wav

-- %USERSIG{DougMiller - 2016-10-18}%
Topic revision: r4 - 01 Nov 2016, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback